# Coupling induced emergent topology in a two-leg fermionic ladder

### v1

@@ Abstract
We investigate the ground state properties of spinless fermions on a two leg ladder, by allowing the nearest-neighbour hopping dimerization in one leg and uniform hopping in the other. In the non-interacting limit, we find that, at half-filling, the system exhibits robust topological behavior if the inter-leg hopping is allowed. As a result, a topological phase transition occurs as a function of the inter-leg hopping. When the inter-leg interaction is turned on, the topological phase survives, and we obtain an interaction induced topological phase transition as a function of interaction. Moreover, we show that for both the non-interacting and interacting cases, topological nature of the system survives even by altering the dimerization pattern in the dimerized leg, resulting in two topological phases of opposite character. Finally, we reveal that when uniform interactions are included on all the bonds of the ladder, the topological phase transitions to a symmetry-broken charge-density wave (CDW) phase.
@@ I Introduction
These ideas have now expanded beyond non-interacting limits, opening new avenues for exploring how strong correlations, dimensionality, and symmetry intertwine to stabilize or destabilize topological order in the free fermionic counterpart Rachel (2018). In this context, SPT serves as the simplest playground to study the interplay of these key aspects of quantum condensed matter physics, as these phases are short-range entangled phases Senthil (2015). Due to this, extensive research has been conducted using interacting SSH models to understand how local and non-local interactions affect topological phases and phase transitions in many-body fermionic, bosonic, and spin systems Grusdt et al. (2013); Di Liberto et al. (2016, 2017); Zhou et al. (2023); Li et al. (2013); Taddia et al. (2017); Fraxanet et al. (2022); Julià-Farré et al. (2024a); Mondal et al. (2021); Melo et al. (2023); Márquez et al. (2024); Zhou et al. (2025). An interesting revelation in this context is the interaction induced topology where interactions alone give rise to nontrivial topology in systems that are otherwise topologically trivial in the non-interacting limit Mondal et al. (2022a); Hetényi (2020); Julià-Farré et al. (2022); Zuo et al. (2020); Parida et al. (2025); Julià-Farré et al. (2024b).
@@
On the other hand, interaction effects and dimensionality of the lattices on the topological phases have been investigated by going beyond strictly one-dimensional chains. In particular, ladder geometries or quasi-one dimensional lattices that are intermediate to one and two-dimensional lattices have attracted a great deal of attention, revealing interesting topological properties  Nersesyan (2020); Zhang and Zhou (2017); Li et al. (2017); Padavić et al. (2018); Padhan et al. (2024); Mondal et al. (2023); Feng et al. (2022); Li et al. (2013); Parida et al. (2024); Downing et al. (2024); Ogino et al. (2021, 2022); Liu et al. (2012). The key component in the ladder systems that plays the major role is the inter-layer coupling that decides the fate of the topological properties of individual chains and also imparts novel physical phenomena due to proximity effects Lu and Wang (2007); Shoman et al. (2015); Hsieh et al. (2016); Cheng et al. (2019); Zheng and Hofstetter (2018).
@@
In this work, we study the topological properties of an asymmetric two-leg ladder model composed of spinless fermions, where one leg follows the SSH-type dimerized hopping, and the other is a uniform tight-binding chain. For the non-interacting case, the system becomes topological for any finite dimerization in the dimerized leg which undergoes a transition to a gapped trivial phase as a function of inter-leg (rung) hopping. Interestingly, in the many-body limit, after including the rung interaction, it also drives a topological phase transition in the system as a function of rung interaction. However, when uniform interactions are introduced along both the legs and rungs, this topological phase transition ceases to happen, and the system instead undergoes a transition into a CDW phase.
@@
The paper is organized as follows. In sec. II, we introduce the model under consideration and the computational method used to obtain all the results. In sec. III, we present the non-interacting case, focusing on both regimes of dimerization in the upper leg. Sec. IV discusses the scenario when finite rung interaction is considered. Finally, in sec. VI, we provide a summary of the key results.
@@ II Model
$\hat{H}$ is a single-particle Hamiltonian of non-interacting spinless fermions. $\hat{a}_{i}(\hat{a}_{i}^{\dagger})$ and $\hat{b}_{i}(\hat{b}_{i}^{\dagger})$ are the particle annihilation (creation) operators on leg-a and leg-b, respectively, where $i$ represents the rung index. $t_{1}$ and $t_{2}$ are the alternating hopping amplitudes in the upper leg, $t$ is the uniform hopping strength in the lower leg, and $t_{p}$ represents the strength of the rung (inter-leg) hopping. Throughout all the numerical calculations, we set $t=1$, which sets the energy scale of the system.
@@
Upon introducing interactions between the fermions, the above model exhibits characteristics of a strongly correlated Fermi system. For simplicity, we consider only rung-wise repulsive interactions, which can be incorporated as the term $\hat{H}_{V}$ in Eq. (1)
$\hat{H}_{V}$ $=V_{p}\sum_{i}(\hat{n}_{i,a}-\frac{1}{2})(\hat{n}_{i,b}-\frac{1}{2}),$ (5)
@@
Figure 2: The left panels in (a-d) represent the single particle dispersion bands for $t_{p}=0.0,~0.8,~\sqrt{2.4},~\text{and}~2.5$ respectively. The right panels in (a-d) represent the corresponding particle filling ($\rho$) plotted against the chemical potential ($\mu$) for the same values of $t_{p}$.
@@ III Results
In this section, we first analyze the effect of rung hopping in the non-interacting ladder, and then we investigate the role of rung interaction.
@@ III.1 Non-interacting case
To understand the topological nature of the non-interacting particles, we begin by examining the gap-opening as well as the gap-closing scenarios in the non-interacting picture. Since in fermionic systems, the energy gaps can be fully understood from the underlying band structure, we start by diagonalizing the Hamiltonian given in Eq. (4), and the resulting dispersion relation is
@@
The above dispersion relation suggests that for any finite dimerization in hopping in leg-a, the gap closes at $k=0$ when $t_{p}^{2}=2t(t_{1}+t_{2})$, and at $k=\pi$, when $t_{p}^{2}=0$. At all other points, the gap remains open. We display the band structure for four exemplary values of $t_{p}$: $t_{p}=0$, $0.8$, $\sqrt{2.4}$, and $2.5$ in the left panles of Fig. 2(a–d) respectively by fixing $t_{1}=0.2$ and $t_{2}=1.0$. To compare the band picture with the filling of particles in real space, the filling $\rho=N/2L$ as a function of chemical potential $\mu$ Crépin et al. (2011) calculated under periodic boundary conditions, is plotted in right panels of Fig. 2(a–d), where $N$ is the total number of fermions and $L$ is the number of rungs. The appearance of cusps in the $\rho$$\mu$ curves reflects changes in the number of Fermi points i.e., the number of quasi-momenta where the Fermi energy intersects the bands. The width of each plateau in the $\rho$$\mu$ curve in Fig. 2 indicates the value of the energy gap at the corresponding filling, whereas the shoulder regions around the plateaus indicate the gapless Luttinger liquid. Comparing the two panels in Fig. 2(a), we observe that for $t_{p}=0$, the system remains gapless at all fillings. In contrast, for $t_{p}=0.8$, finite gaps emerge at quarter, half, and three-quarter fillings, as shown in Fig. 2(b). Increasing the value of $t_{p}$ to $\sqrt{2.4}$ closes the gap at half-filling, while the gaps at quarter and three-quarter fillings persist [see Fig. 2(c)]. At $t_{p}=2.5$, gaps reappear again at all three commensurate fillings, such as for quarter, half, and three-quarter fillings, as shown in Fig. 2(d).
@@
III.2 Topological Phase Transition
@@
Since one of the legs of the ladder is considered to be of SSH type, we can have two possible dimerization patterns: (i) topological pattern ($t_{1}<t_{2}$ as in Fig. 1(a)) and (ii) the trivial pattern ($t_{1}>t_{2}$ as in Fig. 1(b)). For the topological (trivial) pattern, we choose $t_{1}=0.2,~t_{2}=1.0$ ($t_{2}=1.0,~t_{2}=0.2$). Note that the hopping in the other leg is uniform throughout and is set to $t=1.0$. This choice of hopping suggests that in the decoupled leg limit, the leg-a can be either trivial or topological insulator whereas the leg-b is always a gapless Luttinger liquid. Our goal is to investigate the influence of the dimerized leg on the system when the two legs are coupled to each other.
@@
Figure 3: (a) and (b) energy spectrum as a function of rung hopping $t_{p}$ when $t_{1}=0.2$, $t_{2}=1.0$ and $t_{1}=1.0$, $t_{2}=0.2$ respectively. (c) and (d) represent the probability density $|{\psi_{i}}|^{2}$, where the circles represent the sites of the ladder and the facecolor of the circle represents the probability density corresponding to the blue circle in (a) and (b) respectively.
@@
We first explore the real space energy spectrum of the system under OBC. In Fig. 3(a) and (b) we depict the energy spectrum as a function of $t_{p}$ for $t_{1}<t_{2}$ and $t_{1}>t_{2}$ configurations, respectively. For both the configurations, the bulk gaps open up as soon as $t_{p}$ becomes finite. Interestingly, for both the topological and trivial dimerization configuration of leg-a, the bulk gap hosts two mid-gap zero-energy states which disappear at a critical $t_{p}=\sqrt{2.4}$ with the closing up of the bulk gap. However, the gap reopens with further increase in $t_{p}$ without the zero energy states. These zero energy state are typical signatures of topological phases. In order to examine that these zero-energy states are edge states, we plot the onsite probability density ($|{\psi_{i}}|^{2}$) corresponding to one of these two states under OBC. As shown in Fig. 3(c) and (d) the states corresponding to the dimerization patterns $t_{1}=0.2,~t_{2}=1.0$ and $t_{1}=1.0,~t_{2}=0.2$ for an exemplary value $t_{p}=1.0$ (dots in Fig. 3(a) and (b)), are localized at the edges of the system. The key distinction between both the cases is that the edge states exist on the upper leg when $t_{1}<t_{2}$ and on the lower leg when $t_{1}>t_{2}$. Interestingly, we observe the lower-leg, which doesn’t have any hopping dimarization, has supported a zero energy edge state. Additionally, the spectra also exhibit isolated energetic mid-gap states at $\rho=1/4$ and $\rho=3/4$ fillings for $t_{1}<t_{2}$ (the red solid lines), which are absent for $t_{1}>t_{2}$. For half-filling, it is evident that there is a pair of localized edge modes existing in the upper leg for the $t_{1}<t_{2}$ case, whereas at filling $\rho=1/4$, the edges of both the legs of the ladder hybridize, resulting in energetic edge states. In contrast, there is also another pair of edge modes at zero energy for $t_{1}>t_{2}$ case as shown in Fig. 3(d), and these are localized in the lower leg of the ladder. From now on, we focus on the zero energy edge states appearing at half-filling to characterize the topological property of our system, as these are the states protected by symmetry.
@@
where, $k_{j}=2\pi j/Na$ is the $j^{\text{th}}$ $k$ point in the Brillouin zone and $|u_{m,k_{j}}\rangle$ denotes the occupied Bloch bands. The determinant is taken as we are dealing with a multiband system. It is well known that a trivial bulk is characterized by $\omega=0$, while a topological bulk possesses a value $\omega=\pi$. We calculate the Berry phase $\omega$ for both $t_{1}<t_{2}$ as well as $t_{1}>t_{2}$ cases by taking the filled bands up to half filling and plot $\omega/\pi$ as a function of $t_{p}$ in Fig. 4(a) and (b), respectively. At $t_{p}=0$, the two legs of the ladder are completely decoupled, with one forming a tight-binding chain for which the gap closes at $k=\pi$, making the Berry phase undefined. For any finite $t_{p}$, however, the system acquires a well-defined Berry phase, which takes the value $\pi$ for both $t_{1}<t_{2}$ and $t_{1}>t_{2}$ for arbitrarily small $t_{p}$, confirming the topological nature of the phases. As $t_{p}$ increases, the bulk gap closes at $t_{p}=\sqrt{2.4}$, beyond which the Berry phase drops to 0, indicating a transition to a trivial phase. This change in $\omega/\pi$ from $1$ to 0 for both $t_{1}<t_{2}$ and $t_{1}>t_{2}$ provides clear evidence of a topological phase transition at $t_{p}=\sqrt{2.4}$. Furthermore, this transition is also accompanied by the disappearance of the zero-energy edge modes present in the topological phase, and this is called the bulk-boundary correspondence in our system.

### v2

@@ Abstract
We investigate the ground state properties of spinless fermions on a two leg ladder, by allowing the nearest-neighbour hopping dimerization in one leg and uniform hopping in the other. In the non-interacting limit, we find that, at half-filling, the system exhibits robust topological behavior if the inter-leg hopping is allowed. Though depending on the dimerization pattern, the dimerized leg can be either topological or trivial in nature, here we show that by connecting such a leg to a uniform leg through inter-chain coupling, the overall system becomes topological irrespective of the dimerization pattern in the dimerized leg. As a result, a topological phase transition occurs as a function of the inter-leg hopping. When the inter-leg interaction is turned on, the topological phase survives, and we obtain an interaction induced topological phase transition. Finally, we reveal that when uniform interactions are included on all the bonds of the ladder, the topological phase transitions to a symmetry-broken charge-density wave (CDW) phase.
@@ I Introduction
These ideas have now expanded beyond non-interacting limits, opening new avenues for exploring how strong correlations, dimensionality, and symmetry intertwine to stabilize or destabilize topological order in the free fermionic counterpart Rachel (2018). In this context, SPT serves as the simplest playground to study the interplay of these key aspects of quantum condensed matter physics, as these phases are short-range entangled phases Chen et al. (2010); Senthil (2015). Due to this, extensive research has been conducted using interacting SSH models to understand how local and non-local interactions affect topological phases and phase transitions in many-body fermionic, bosonic, and spin systems Grusdt et al. (2013); Di Liberto et al. (2016, 2017); Zhou et al. (2023); Li et al. (2013); Taddia et al. (2017); Fraxanet et al. (2022); Julià-Farré et al. (2024a); Mondal et al. (2021); Melo et al. (2023); Márquez et al. (2024); Zhou et al. (2025); Mohamadi and Abouie (2025). An interesting revelation in this context is the interaction induced topology where interactions alone give rise to nontrivial topology in systems that are otherwise topologically trivial in the non-interacting limit Mondal et al. (2022a); Hetényi (2020); Julià-Farré et al. (2022); Zuo et al. (2020); Parida et al. (2025); Julià-Farré et al. (2024b).
@@
On the other hand, interaction effects and dimensionality of the lattices on the topological phases have been investigated by going beyond strictly one-dimensional chains. In particular, ladder geometries or quasi-one dimensional lattices that are intermediate to one and two-dimensional lattices have attracted a great deal of attention, revealing interesting topological properties  Nersesyan (2020); Zhang and Zhou (2017); Li et al. (2017); Padavić et al. (2018); Padhan et al. (2024); Mondal et al. (2023); Feng et al. (2022); Li et al. (2013); Parida et al. (2024); Downing et al. (2024); Ogino et al. (2021, 2022); Liu et al. (2012); Aghtouman and Hosseini (2024); D’Elia and Paiva (2025); Jangjan and Hosseini (2020, 2022). The key component in the ladder systems that plays the major role is the inter-layer coupling that decides the fate of the topological properties of individual chains and may impart novel physical phenomena due to proximity effects Lu and Wang (2007); Shoman et al. (2015); Hsieh et al. (2016); Cheng et al. (2019); Zheng and Hofstetter (2018).
@@
In this work, we study the topological properties of an asymmetric two-leg ladder model composed of spinless fermions, where one leg follows the SSH-type dimerized hopping, and the other is a uniform tight-binding chain. For the non-interacting case, the dimerization pattern decides the topological or trivial nature of the SSH-type leg. However, once coupled to a uniform tight-binding chain through interchain tunneling, the ladder as a whole becomes topological regardless of the dimerization pattern in the dimerized leg. Interestingly, in the many-body limit, including the rung interaction, it also drives a topological phase transition in the system as a function of rung interaction. However, when uniform interactions are introduced along both the legs and rungs, this topological phase transition ceases to happen, and the system instead undergoes a transition into a CDW phase.
@@
The paper is organized as follows. In sec. II, we introduce the model under consideration and the computational method used to obtain all the results. In sec. III.1, we present the non-interacting case, focusing on both regimes of dimerization in the SSH-type leg. Sec. III.2 discusses the scenario when interaction between the particles is considered. Finally, in sec. IV, we provide a summary of the key results.
@@ II Model
$\hat{H}$ represents the Hamiltonian of non-interacting spinless fermions. $\hat{a}_{i}(\hat{a}_{i}^{\dagger})$ and $\hat{b}_{i}(\hat{b}_{i}^{\dagger})$ are the particle annihilation (creation) operators on leg-a and leg-b, respectively, where $i$ represents the rung index. $t_{1}$ and $t_{2}$ are the alternating hopping amplitudes in the upper leg, $t$ is the uniform hopping strength in the lower leg, and $t_{p}$ represents the strength of the rung (inter-leg) hopping. Throughout all the numerical calculations, we set $t=1$, which sets the energy scale of the system.
@@
Upon introducing interactions between the fermions, the above model exhibits characteristics of a strongly correlated Fermi system. For simplicity, we consider only rung-wise repulsive interactions, which can be incorporated as the term $\hat{H}_{V_{p}}$ in Eq. (1)
$\hat{H}_{V_{p}}$ $=V_{p}\sum_{i}(\hat{n}_{i,a}-\frac{1}{2})(\hat{n}_{i,b}-\frac{1}{2}),$ (5)
@@
Figure 2: The left panels in (a-d) represent the single particle dispersion bands for $t_{p}=0.0,~0.8,~\sqrt{2.4},~\text{and}~2.5$ respectively. The right panels in (a-d) represent the corresponding particle filling ($\rho$) plotted against the chemical potential ($\mu$) for the same values of $t_{p}$ with a system of $L=500$ rungs. For all the plots, the values of $t_{1}$ and $t_{2}$ are fixed at $0.2$ and $1.0$ respectively.
@@ III Results
In this section, we first analyze the noninteracting case, where we investigate the effect of rung hopping, and then we investigate the role of interparticle interaction.
@@ III.1 Non-interacting case
To understand the topological nature of the non-interacting particles, we begin by examining the gap-opening as well as the gap-closing scenarios. Since in fermionic systems, the energy gaps can be fully understood from the underlying band structure, we start by diagonalizing the Hamiltonian given in Eq. (4), and the resulting dispersion relation is
@@
The above dispersion relation suggests that for any finite dimerization in hopping in leg-a, the gap closes at $k=0$ when $t_{p}^{2}=2t(t_{1}+t_{2})$, and at $k=\pi$, when $t_{p}^{2}=0$. At all other points, the gap remains open at half-filling. We display the band structure for four exemplary values of $t_{p}$: $t_{p}=0$, $0.8$, $\sqrt{2.4}$, and $2.5$ in the left panles of Fig. 2(a–d) respectively by fixing $t_{1}=0.2$ and $t_{2}=1.0$. To compare the band picture with the filling of particles in real space, the filling $\rho=N/2L$ as a function of chemical potential $\mu$ Crépin et al. (2011) calculated under PBC, is plotted in right panels of Fig. 2(a–d), where $N$ is the total number of fermions and $L$ is the number of rungs. The appearance of cusps in the $\rho$$\mu$ curves reflects changes in the number of Fermi points i.e., the number of quasi-momenta where the Fermi energy intersects the bands. The width of each plateau in the $\rho$$\mu$ curve indicates the value of the energy gap at the corresponding filling, whereas the shoulder regions around the plateaus indicate the gapless Luttinger liquid. Comparing the two panels in Fig. 2(a), we observe that for $t_{p}=0$, the system remains gapless at all fillings. In contrast, for $t_{p}=0.8$, finite gaps emerge at $\rho=1/4,~1/2,~\text{and}~3/4$, as shown in Fig. 2(b). Increasing the value of $t_{p}$ to $\sqrt{2.4}$ closes the gap at $\rho=1/2$, while the gaps at $\rho=1/4,~\text{and}~3/4$ persist [see Fig. 2(c)]. At $t_{p}=2.5$, gaps reappear again at all three commensurate fillings, such as for $\rho=1/4,~1/2,~\text{and}~3/4$, as shown in Fig. 2(d).
@@
III.1.1 Topological Phase Transition
@@
Since one of the legs of the ladder is considered to be of SSH type, we can have two possible dimerization patterns: (i) topological pattern ($t_{1}<t_{2}$ as in Fig. 1(a)) and (ii) the trivial pattern ($t_{1}>t_{2}$ as in Fig. 1(b)). For the topological (trivial) pattern, we choose $t_{1}=0.2,~t_{2}=1.0$ ($t_{2}=1.0,~t_{2}=0.2$). Note that the hopping in the other leg is uniform throughout and is set to $t=1.0$. This choice of hopping suggests that in the decoupled leg limit, the leg-a can be either trivial or a topological insulator whereas the leg-b is always a gapless Luttinger liquid. Our goal is to investigate the influence of the dimerized leg on the system when the two legs are coupled to each other.
@@
Figure 3: (a) and (b) energy spectrum as a function of rung hopping $t_{p}$ when $t_{1}=0.2$, $t_{2}=1.0$ and $t_{1}=1.0$, $t_{2}=0.2$ respectively. (c) and (d) represent the probability density $|{\psi_{i}}|^{2}$, where the circles represent the sites of the ladder and the facecolor of the circle represents the probability density corresponding to the blue circle in (a) and (b), respectively.
@@
We first explore the real space energy spectrum of the system under OBC. In Fig. 3(a) and (b) we depict the energy spectrum as a function of $t_{p}$ for $t_{1}<t_{2}$ and $t_{1}>t_{2}$ configurations, respectively. For both the configurations, the bulk gaps open up as soon as $t_{p}$ becomes finite. Interestingly, for both the topological and trivial dimerization configuration of leg-a, the bulk gap hosts two mid-gap zero-energy states which disappear at a critical $t_{p}=\sqrt{2.4}$ with the closing up of the bulk gap. However, the gap reopens with further increase in $t_{p}$ without the zero energy states. These zero energy states are typical signatures of topological phases under OBC. In order to examine that these zero-energy states are edge states, we plot the onsite probability density ($|{\psi_{i}}|^{2}$) corresponding to one of these two states. As shown in Fig. 3(c) and (d) the states corresponding to the dimerization patterns $t_{1}=0.2,~t_{2}=1.0$ and $t_{1}=1.0,~t_{2}=0.2$ for an exemplary value $t_{p}=1.0$ (dots in Fig. 3(a) and (b)), are localized at the edges of the system. The key distinction between both the cases is that the edge states exist on the upper leg when $t_{1}<t_{2}$ and on the lower leg when $t_{1}>t_{2}$. Interestingly, we observe the lower-leg, which doesn’t have any hopping dimarization, has supported a zero energy edge state. Additionally, the spectra also exhibit isolated energetic mid-gap states at $\rho=1/4$ and $\rho=3/4$ fillings for $t_{1}<t_{2}$ (the red solid lines), which are absent for $t_{1}>t_{2}$. For $\rho=1/2$, it is evident that there is a pair of localized edge modes existing in the upper leg for the $t_{1}<t_{2}$ case, whereas at $\rho=1/4$, the edges of both the legs of the ladder hybridize (not shown), resulting in energetic edge states. In contrast, there is also another pair of edge modes at zero energy for $t_{1}>t_{2}$ case as shown in Fig. 3(d), and these are localized in the lower leg of the ladder. From now on, we focus on the zero energy edge states appearing at half-filling to characterize the topological property of our system, as these are the states protected by symmetry.
@@
where, $\mathfrak{I}$ denotes the imaginary part, $k_{j}=2\pi j/Na$ is the $j^{\text{th}}$ $k$ point in the Brillouin zone and $|u_{m,k_{j}}\rangle$ denotes the occupied Bloch bands with $m$ and $n$ representing the band indices. The determinant is taken as we are dealing with a multiband system. It is well known that a trivial bulk is characterized by $\omega=0$, while a topological bulk possesses a value $\omega=\pi$. We calculate the Berry phase $\omega$ for both $t_{1}<t_{2}$ as well as $t_{1}>t_{2}$ cases by taking the filled bands up to half filling and plot $\omega/\pi$, called the winding number, as a function of $t_{p}$ in Fig. 4(a) and (b), respectively. At $t_{p}=0$, the two legs of the ladder are completely decoupled, with one forming a tight-binding chain for which the gap closes at $k=\pi$, making the Berry phase undefined. For any finite $t_{p}$, however, the system acquires a well-defined Berry phase, which takes the value $\pi$ for both $t_{1}<t_{2}$ and $t_{1}>t_{2}$ for arbitrarily small $t_{p}$, confirming the topological nature of the phases. As $t_{p}$ increases, the bulk gap closes at $t_{p}=\sqrt{2.4}$, beyond which the Berry phase drops to 0, indicating a transition to a trivial phase. This change in $\omega/\pi$ from $1$ to 0 for both $t_{1}<t_{2}$ and $t_{1}>t_{2}$ provides clear evidence of a topological phase transition at $t_{p}=\sqrt{2.4}$. Furthermore, this transition is also accompanied by the disappearance of the zero-energy edge modes present in the topological phase, and this is called the bulk-boundary correspondence in our system.
III.1.2 Phase diagram